OpenStudy (anonymous):

Put the following into a function form 5x+15y=30

OpenStudy (anonymous):

you want to annoy your teacher? put the totally correct, but not expected answer of \[f:=\{(x,y)|5x+15y=30\}\]

OpenStudy (anonymous):

what they want you to do is solve this equation for \(y\) which takes only two steps

OpenStudy (anonymous):

subtract \(5x\) from both sides to get \[15y=30-5x\] then divide both sides by \(15\) and get \[y=2-\frac{1}{3}x\] or \[f(x)=2-\frac{1}{3}x\]

OpenStudy (anonymous):

i still think you should write \[\large f:=\{(x,y)|5x+15y=30\}\] since that is the real right answer

OpenStudy (anonymous):

thank you